OpenStudy (anonymous):

patricia wants to build a mansion. the perimeter of the land she need is at least, 188m. the length of the property has to be 34 m. more than the width. Find the width and length of the land

OpenStudy (nowhereman):

Name the width as a variable x and write down the given property in terms of x.

OpenStudy (anonymous):

If we set x as the width and y as the length the perimeter is given by 2x + 2y = 188 km. We also know x + 34 = y. If we substitute this in to our first equation we get: 2x + 2(x + 34) = 188. Expanding 2x + 2x + 68 = 188. 4x = 120 so x = 30 and y = 64.

OpenStudy (anonymous):

circumference = 2xL + 2xW ==> 188 = 2xL + 2xW also W = L + 34 substituting W into the first equation ==> 188 = 2xL + 2x(L+24) ==> solve for L ==> L = 30 subsitiute L into original or second equations to solve for W ==> W = 30+34 = 64